# BiLO: Bilevel Local Operator Learning for PDE Inverse Problems. Part II: Efficient Uncertainty Quantification with Low-Rank Adaptation

### v1

@@ Abstract
Uncertainty quantification and inverse problems governed by partial differential equations (PDEs) are central to a wide range of scientific and engineering applications. In this second part of a two part series, we extend Bilevel Local Operator Learning (BiLO) for PDE-constrained optimization problems developed in Part 1 to the Bayesian inference framework. At the lower level, we train a network to approximate the local solution operator by minimizing the local operator loss with respect to the weights of the neural network. At the upper level, we sample the PDE parameters from the posterior distribution. We achieve efficient sampling through gradient-based Markov Chain Monte Carlo (MCMC) methods and low-rank adaptation (LoRA). Compared with existing methods based on Bayesian neural networks, our approach bypasses the challenge of sampling in the high-dimensional space of neural network weights and does not require specifying a prior distribution on the neural network solution. Instead, uncertainty propagates naturally from the data through the PDE constraints. By enforcing strong PDE constraints, the proposed method improves the accuracy of both parameter inference and uncertainty quantification. We analyze the dynamic error of the gradient in the MCMC sampler and the static error in the posterior distribution due to inexact minimization of the lower level problem and demonstrate a direct link between the tolerance for solving the lower level problem and the accuracy of the resulting uncertainty quantification. Through numerical experiments across a variety of PDE models, we demonstrate that our method delivers accurate inference and quantification of uncertainties while maintaining high computational efficiency.
@@ 1 Introduction
Quantifying uncertainty in physical systems is crucial for informed decision-making across various disciplines. Key applications include seismic imaging [1, 2, 3, 4], personalized medicine [5, 6], climate modeling [7], and engineering design [8]. These fields frequently require solving PDE inverse problems within a Bayesian framework [9], where the objective is to estimate the posterior distribution of model parameters given observed data. Typically, this process involves sampling PDE parameters and repeatedly solving the forward PDE. A related approach, PDE-constrained optimization, aims to identify a single optimal parameter estimate, thus offering computational efficiency but lacking the ability to quantify uncertainty.
This is the second paper in a two-part series. In Part I, we develop a constrained optimization framework to solve inverse PDE problems using deep learning methods [10]. Here in Part II, we extend this approach to Bayesian inference frameworks building upon the core methodology developed in Part I, while leveraging transfer learning techniques to enhance sampling efficiency.
@@
Addressing Bayesian PDE inverse problems involves two design choices: the PDE solver and the sampling method. PDE solutions can be computed using numerical methods (such as finite difference or finite element methods), physics-informed neural networks (PINNs)[11], or neural operators such as Fourier Neural Operator [12] and DeepONet [13]. Each approach has its advantages and limitations. Numerical methods provide fast, accurate, and convergent solutions but can struggle with high-dimensional PDEs [14]. PINNs offer a mesh-free approach, effective in some higher-dimensional PDE problems, and a unified framework for inverse problems. Neural operators, while requiring many synthetic numerical solutions to train, enable rapid evaluations and serve as PDE surrogates; however, their accuracy remains relatively low, especially outside the distribution of the training dataset. For sampling, Markov Chain Monte Carlo (MCMC) [15] methods can be either gradient-based (e.g., Hamiltonian Monte Carlo [16, 17, 18]) or derivative-free (e.g., Ensemble Kalman Sampling [19]). Gradient-based methods efficiently explore posterior distributions, but computing the gradient can be expensive (compared with gradient-free approaches) or impractical (e.g., when the PDE solver is a black box).
@@
Numerical PDE solvers can be paired with gradient-based sampling methods. The Adjoint Method computes gradients of the likelihood function with respect to PDE parameters [20]. While the adjoint method ensures accurate solutions and gradients, deriving adjoint equations for complex PDEs can be challenging and solving both the forward and adjoint equations numerically can be computationally intensive. When only the forward solver is available, it is more practical to use derivative-free sampling methods such as Ensemble Kalman Sampling (EnKS), which employ ensembles of interacting particles to approximate posterior distributions [19]. Additionally, variants of Ensemble Kalman Inversion (EnKI) can efficiently estimate the first two moments in Bayesian settings [21, 22, 23].
@@
Among neural PDE solvers, Physics-Informed Neural Networks (PINNs) [11] have become widely adopted. In PINNs, the PDE residual is enforced as a soft constraint. By minimizing a weighted sum of the data loss and the PDE residual loss, PINNs can fit the data, solve the PDE, and infer the parameters at the same time. For uncertainty quantification, Bayesian PINNs (BPINNs) [24] represent PDE solutions using Bayesian neural networks, and use Hamiltonian Monte Carlo (HMC) to sample the joint posterior distributions of PDE parameters and neural network weights. Derivative-free methods, such as EnKI, can also be effectively combined with neural PDE solvers, as demonstrated in PDE-constrained optimization problems [25] and Bayesian PINNs [23].
@@
Neural Operators (NOs) approximate the PDE solution operator (parameter-to-solution map), providing efficient surrogate models for forward PDE solvers [26]. Once trained, these surrogates integrate seamlessly with Bayesian inference frameworks, leveraging rapid neural network evaluations for posterior distribution sampling using both gradient-based and derivative-free methods [27, 28, 29]. Prominent examples include the Fourier Neural Operator (FNO) [12, 30, 31, 32], Deep Operator Network (DeepONet) [13, 33], and In-Context Operator (ICON) [34], among others [35, 36]. Nevertheless, neural operator performance heavily depends on the training dataset, potentially limiting accuracy in PDE solutions and consequently degrading posterior distribution fidelity [37].
In this work, we focus explicitly on addressing PDE inverse problems using the Bayesian framework through neural PDE solvers and gradient-based MCMC methods. The key contributions of this work are as follows:
Main Contributions
@@
  • We introduce a Bilevel local operator (BiLO) learning approach that enforces strong PDE constraints, leading to improved accuracy in uncertainty quantification and parameter inference.
@@
  • We apply Low-Rank Adaptation (LoRA)[38] to enhance both efficiency and speed in fine-tuning and sampling, significantly reducing computational cost while maintaining accuracy.
@@
The outline of the paper is as follows. In Sec. 2, we present the BiLO framework for Bayesian inference with low rank adaption (LoRA). We highlight the differences between our approach and those used in BPINNs. In Sec. 3, we present numerical results for a variety of PDE inverse problems, and compare the results with BPINNs. In Sec. 4, we summarize the results. In the Appendices, we present theoretical analyses and additional numerical results.
2 Method
In this section, we present the BiLO framework for Bayesian inference with low-rank adaptation (LoRA). The section is organized as follows. In 2.1 we formulate the PDE-based Bayesian inverse problem as a bi-level optimization problem, where the upper level involves sampling the PDE parameters, and the lower level trains a neural network to approximate the local operator. In 2.2, we describe the neural network architecture used to approximate the local operator. In 2.3, we provide a brief overview of the Hamiltonian Monte Carlo (HMC) algorithm in a general context, independent of the BiLO framework. In 2.4, we present our main algorithm—BiLevel Local Operator Hamiltonian Monte Carlo (BiLO-HMC)—along with several theoretical results. In 2.5, we describe the pre-training and fine-tuning workflow, and introduce the low-rank adaptation (LoRA) technique used to accelerate inference. In 2.6, we highlight the differences between our approach and Bayesian physics-informed neural networks (BPINNs) [24].
2.1 Bi-Level Local Operator for Bayesian Inference
Let $u:\Omega\to\mathbb{R}$ be a function defined over a domain $\Omega\subset\mathbb{R}^{d}$ satisfying some boundary conditions. Suppose $u$ is governed by some constraints that include the PDE, boundary conditions, and initial conditions,
$\mathcal{F}(D^{k}u(\mathbf{x}),...,Du(\mathbf{x}),u(\mathbf{x}),\theta)=\mathbf{0},$ (1)
where $D^{k}$ is the $k$-th order derivative operator and $\theta$ are the PDE parameters. We assume that the solution is unique for each $\theta$ and that the solution depends smoothly on $\theta$. The precise assumptions on the function space will be specified when a theorem is presented.
We consider functions of the form $u(\mathbf{x},\theta)$ and define the residual function
$r(\mathbf{x},\theta):=\mathcal{F}(D^{k}u(\mathbf{x},\theta),...,Du(\mathbf{x},\theta),u(\mathbf{x},\theta),\theta)\$ (2)
The dependence of $r$ on $u$ is not explicitly indicated for brevity. A function $u(\mathbf{x},\theta)$ is called a full operator (hereafter referred to as the “operator”) if $r(\mathbf{x},\theta)=0$ for all $\theta$; that is, it solves the PDE for all $\theta$ in some admissible set. We define $u$ as a local operator at $\theta$ if the following two conditions hold:
  • Condition 1: $r(\mathbf{x},\theta)=\mathbf{0}$.
  • Condition 2: $\gradient_{\theta}r(\mathbf{x},\theta)=d_{\theta}\mathcal{F}(D^{k}u(\mathbf{x},\theta),...,Du(\mathbf{x},\theta),u(\mathbf{x},\theta),\theta)=\mathbf{0}$.
where $d_{\theta}$ is the total derivative with respect to $\theta$.
@@ 2.1 Bi-Level Local Operator for Bayesian Inference
For Bayesian inference, let $P(\theta)$ be the prior distribution of the PDE parameters, let $D$ be the observed data, $P(D|\theta)$ be the likelihood of the observed data given the PDE parameters, and $P(\theta|D)$ be the posterior distribution of the PDE parameters. For example, if we assume that the data is observed at $N$ collocation points $\mathbf{x}_{i}\in\Omega$, and the observation is corrupted by Gaussian noise with mean 0 and variance $\sigma_{d}^{2}$, then the likelihood of the observed data $D$ given the PDE parameters $\theta$ is
$P(D|\theta)=\prod_{i=1}^{N}\frac{1}{\sqrt{2\pi\sigma_{d}^{2}}}\exp\left(-\frac{1}{2\sigma_{d}^{2}}\left|u(\mathbf{x}_{i},\theta)-\hat{u}(\mathbf{x}_{i})\right|^{2}\right)$ (3)
@@
By Bayes’ theorem, $P(\theta|D)\propto P(D|\theta)P(\theta)$. The potential energy, $U(\theta)$, is defined as the negative of the log-posterior distribution
$U(\theta)=-\log P(\theta|D)=-\log P(D|\theta)-\log P(\theta)$ (4)
@@
The potential energy $U$ is a functional of the PDE solution $u(\cdot,\theta)$, and thus ultimately depends only on the PDE parameters $\theta$. Sampling the posterior distribution with gradient-based MCMCs, such as Langevin Dynamics (LD) or Hamiltonian Monte Carlo (HMC), requires computing the gradient of the potential energy $\nabla_{\theta}U(\theta)$ to effectively explore the posterior distribution.
@@
In this work, we represent the local operator by a neural network $u(\mathbf{x},\theta;W)$, where $W$ are the weights of the neural network. Then $u(\cdot,\theta;W)$ is the approximate solution of the PDE at $\theta$. The detailed architecture of the neural network will be described in Section 2.2. The residual function depends on the weights of the neural network $W$, and can be written as $r(\mathbf{x},\theta;W)$. The two conditions for $u$ to be a local operator at $\theta$ lead to the definition of the local operator loss $\mathcal{L}_{\text{LO}}$:
$\mathcal{L}_{\text{LO}}(\theta,W)=\norm{r(\cdot,\theta;W)}^{2}_{2}+w_{\rm rgrad}\norm{\nabla_{\theta}r(\cdot,\theta;W)}^{2}_{2},$ (5)
@@
The 2-norm of the residual and its gradient with respect to $\theta$ are approximated by the Mean Squared Error (MSE) of the residual and the residual-gradient loss at a set of collocation points. The weight $w_{\rm rgrad}>0$ is a scalar hyperparameter that controls the relative importance of the residual gradient loss. In theory, both terms can be sufficiently minimized to near zero. For fixed $\theta$, by minimizing the local operator loss $\mathcal{L}_{\text{LO}}(\theta,W)$ with respect to the weights $W$, we approximate the local operator $u(\mathbf{x},\theta;W)$ around $\theta$. Since $U$ depends on the approximate solution of the PDE $u(\cdot,\theta;W)$ at $\theta$, we can also write the potential energy as $U(\theta,W)$.
@@
We consider the following Bayesian inference problem:
$\begin{cases}\theta\sim\exp(-U(\theta,W^{*}(\theta)))\\ W^{*}(\theta)=\arg\min_{W}\mathcal{L}_{\text{LO}}(\theta,W)\\ \end{cases}$ (6)
@@
At the lower level, we train a network to approximate the local operator $u(\mathbf{x},\theta;W)$ by minimizing the local operator loss with respect to the weights of the neural network. At the upper level, we sample the PDE parameters $\theta$ from the posterior distribution $P(\theta|D)$.
Hereafter, for notational convenience, we assume that quantities that do not depend explicitly on $W$ are exact. For example, $u(\mathbf{x},\theta)$ is used to denote the exact local or global operator (depending on context), and $u(\cdot,\theta)$ is the exact solution of the PDE at $\theta$. PDE solutions computed using numerical methods (e.g., finite element or finite difference methods) on a fine grid are also considered exact, as they are generally more accurate than neural network approximations [14, 39]. $U(\theta)$ denotes the potential energy of the exact PDE solution $u(\cdot,\theta)$ given $\theta$. The explicit presence of $W$ is taken to indicate neural network approximations of the corresponding quantities. For example, $u(\mathbf{x},\theta;W)$ is the neural network approximation of the local operator at $\theta$, and $U(\theta,W)$ denotes the potential energy of the function $u(\cdot,\theta;W)$.
A schematic of the BiLO framework is shown in Figure 1. The figure is based on a model boundary value problem $-\theta u_{xx}=\sin(\pi x)$ with the Dirichlet boundary condition $u(0)=u(1)=0$. In this case, the full operator is given by $u(x,\theta)=\sin(\pi x)/(\pi^{2}\theta)$, which solves the PDE for all $\theta>0$.
@@
Figure 1 panel (A) illustrates the upper-level problem: the goal is to sample the PDE parameters $\theta$ from the posterior distribution. For gradient-based MCMC methods, we need to compute the gradient of the potential energy $\nabla_{\theta}U(\theta)$ at various values of $\theta$ (orange arrows). Therefore, at any fixed $\theta$, we need access to the local behavior of the potential energy, shown as the blue strip near $\theta_{1}$ and $\theta_{2}$.
@@
In Figure 1 panel (B), the gray surface is the full PDE operator $u(x,\theta)$. At some particular $\theta_{1}$, the local operator $u(x,\theta_{1})$ approximates a small neighborhood of the full operator near $\theta_{1}$ (blue surface). The condition of the local operator is that the residual and the gradient of the residual with respect to $\theta$ are both zero at any fixed $\theta$. Knowing the local operator at $\theta_{1}$ allows us to compute the gradient of the potential energy at $\theta_{1}$.

### v2

# Bayesian BiLO: Bilevel Local Operator Learning for Efficient Uncertainty Quantification of Bayesian PDE Inverse Problems with Low-Rank Adaptation

@@ Abstract
Uncertainty quantification in PDE inverse problems is essential in many applications. Scientific machine learning and AI enable data-driven learning of model components while preserving physical structure, and provide the scalability and adaptability needed for emerging imaging technologies and clinical insights. We develop a Bilevel Local Operator Learning framework for Bayesian inference in PDEs (B-BiLO). At the upper level, we sample parameters from the posterior via Hamiltonian Monte Carlo, while at the lower level we fine-tune a neural network via low-rank adaptation (LoRA) to approximate the solution operator locally. B-BiLO enables efficient gradient-based sampling without synthetic data or adjoint equations and avoids sampling in high-dimensional weight space, as in Bayesian neural networks, by optimizing weights deterministically. We analyze errors from approximate lower-level optimization and establish their impact on posterior accuracy. Numerical experiments across PDE models, including tumor growth, demonstrate that B-BiLO achieves accurate and efficient uncertainty quantification.
@@ 1 Introduction
Uncertainty quantification in computational models governed by partial differential equations (PDEs) is essential across scientific and engineering disciplines, including seismic imaging [dengOpenFWILargeScaleMultiStructural2023, martinStochasticNewtonMCMC2012, yangSeismicWavePropagation2021, bui-thanhExtremescaleUQBayesian2012], personalized medicine [lipkovaPersonalizedRadiotherapyDesign2019, chaudhuriPredictiveDigitalTwin2023], climate modeling [senGlobalOptimizationMethods2013], and engineering design [cookOptimizationTurbulenceModel2019]. These fields frequently require solving PDE inverse problems within a Bayesian framework [stuartInverseProblemsBayesian2010], where the objective is to estimate the posterior distribution of model parameters given observed data. Typically, this process involves sampling PDE parameters and repeatedly solving the forward PDE. A related approach, PDE-constrained optimization, aims to identify a single optimal parameter estimate, thus offering computational efficiency but lacking the ability to quantify uncertainty. Beyond the cost and challenges of solving PDEs, mathematical models may be misspecified due to limited scientific understanding and therefore require continual refinement as new data become available. Scientific machine learning offers a framework for learning components of PDE models directly from data while retaining physical structure. Moreover, the rapid pace of new imaging technologies and clinical insights demands computational tools that are fast, scalable, and adaptable—capabilities well aligned with modern AI/ML methods. Such needs have driven growing interest in scientific machine learning methods that incorporate data-driven learning of model components.
The Bilevel Local Operator learning (BiLO) method has recently shown promise for solving PDE inverse problems in the constrained optimization framework [zhangBiLOBilevelLocal2025a]. A brief review of BiLO is provided in the supplementary material (SM). In this work, we introduce a novel framework, Bayesian BiLO (B-BiLO), which redesigns this bilevel structure for the more challenging setting of Bayesian inference. B-BiLO enables robust uncertainty quantification by integrating local operator learning with gradient-based MCMC sampling and transfer learning techniques for efficiency.
@@
Addressing Bayesian PDE inverse problems involves two design choices: the PDE solver and the sampling method. PDE solutions can be computed using numerical methods (such as finite difference or finite element methods), physics-informed neural networks (PINNs)[raissiPhysicsinformedNeuralNetworks2019], or neural operators such as Fourier Neural Operator [liFourierNeuralOperator2021] and DeepONet [luLearningNonlinearOperators2021]. Each approach has its advantages and limitations. Numerical methods provide fast, accurate, and convergent solutions but can struggle with high-dimensional PDEs [grossmannCanPhysicsinformedNeural2024]. PINNs offer a mesh-free approach, effective in some higher-dimensional PDE problems, and a unified framework for inverse problems. Neural operators, while requiring many synthetic numerical solutions to train, enable rapid evaluations and serve as PDE surrogates; however, their accuracy typically degrades when evaluated outside the distribution of the training dataset [nelsenOperatorLearningMeets2025]. For sampling, Markov Chain Monte Carlo (MCMC) [hastingsMonteCarloSampling1970] methods can be either gradient-based (e.g., Hamiltonian Monte Carlo [nealMCMCUsingHamiltonian2011, nealBayesianLearningNeural1996, girolamiRiemannManifoldLangevin2011]) or derivative-free (e.g., Ensemble Kalman Sampling [garbuno-inigoInteractingLangevinDiffusions2020]). Gradient-based methods efficiently explore posterior distributions, but computing the gradient can be expensive (compared with gradient-free approaches) or impractical (e.g., when the PDE solver is a black box).
@@
Numerical PDE solvers can be paired with gradient-based sampling methods. The Adjoint Method computes gradients of the likelihood function with respect to PDE parameters [bui-thanhSolvingLargescalePDEconstrained2014]. While the adjoint method ensures accurate solutions and gradients, deriving and regularizing the adjoint equations for complex PDEs can be challenging and solving both the forward and adjoint equations numerically can be computationally intensive. When only the forward solver is available, it is more practical to use derivative-free sampling methods such as Ensemble Kalman Sampling (EnKS), which employ ensembles of interacting particles to approximate posterior distributions [garbuno-inigoInteractingLangevinDiffusions2020]. Additionally, variants of Ensemble Kalman Inversion (EnKI) can efficiently estimate the first two moments in Bayesian settings [huangEfficientDerivativefreeBayesian2022, huangIteratedKalmanMethodology2022, pensoneaultEfficientBayesianPhysics2024].
@@
Among neural PDE solvers, Physics-Informed Neural Networks (PINNs) [raissiPhysicsinformedNeuralNetworks2019] have become widely adopted. In PINNs, the PDE residual is enforced as a soft constraint. By minimizing a weighted sum of the data loss and the PDE residual loss, PINNs can fit the data, solve the PDE, and infer the parameters at the same time. For uncertainty quantification, Bayesian PINNs (BPINNs) [yangBPINNsBayesianPhysicsinformed2021] represent PDE solutions using Bayesian neural networks, and use Hamiltonian Monte Carlo (HMC) [duaneHybridMonteCarlo1987, nealBayesianLearningNeural1996] to sample the joint posterior distributions of PDE parameters and neural network weights. Derivative-free methods, such as EnKI, can also be effectively combined with neural PDE solvers, as demonstrated in PDE-constrained optimization problems [guthEnsembleKalmanFilter2020] and Bayesian PINNs [pensoneaultEfficientBayesianPhysics2024].
@@
Neural Operators (NOs) approximate the PDE solution operator (parameter-to-solution map), providing efficient surrogate models for forward PDE solvers [kovachkiNeuralOperatorLearning2022]. Once trained, these surrogates integrate seamlessly with Bayesian inference frameworks, leveraging rapid neural network evaluations for posterior distribution sampling using both gradient-based and derivative-free methods [pathakFourCastNetGlobalDatadriven2022, luMultifidelityDeepNeural2022, maoPPDONetDeepOperator2023]. Prominent examples include the Fourier Neural Operator (FNO) [liFourierNeuralOperator2021, liPhysicsInformedNeuralOperator2024, whitePhysicsInformedNeuralOperators2023, akyildizEfficientPriorCalibration2025], Deep Operator Network (DeepONet) [luLearningNonlinearOperators2021, wangLearningSolutionOperator2021], and In-Context Operator (ICON) [yangIncontextOperatorLearning2023], among others [oleary-roseberryDerivativeInformedNeuralOperator2024, molinaroNeuralInverseOperators2023]. Nevertheless, neural operator performance heavily depends on the training dataset, potentially limiting accuracy in PDE solutions and consequently degrading posterior distribution fidelity [caoResidualbasedErrorCorrection2023].
Our framework integrates key practical advantages from existing sampling and PDE-solving methods. We employ gradient-based MCMC, which generally explores the posterior distribution more efficiently than derivative-free approaches. The PDE solution and the corresponding likelihood gradients with respect to parameters are computed using a neural solver within a novel bilevel optimization framework. This design is flexible across PDE models, and does not require either deriving the adjoint equations or curating pretraining datasets. Our methodology improves accuracy through strong enforcement of PDE constraints without balancing data fitting and PDE residuals as in PINNs, and enhances efficiency by optimizing network weights deterministically rather than sampling in a high-dimensional weight space as in Bayesian neural networks. The key contributions of this work are as follows:
@@
  • We introduce a Bayesian Bilevel local operator (B-BiLO) learning approach that enforces strong PDE constraints, leading to improved accuracy in uncertainty quantification and parameter inference.
@@
  • We apply Low-Rank Adaptation (LoRA)[huLoRALowRankAdaptation2021] to enhance both efficiency and speed in fine-tuning and sampling, significantly reducing computational cost while maintaining accuracy.
@@
2 Results
2.1 Method Overview
This section provides an overview of B-BILO. Full methodological details and theoretical analysis are provided in Section 4 and the Supplementary Material (SM).
PDE Bayesian Inverse Problem
We consider problems governed by partial differential equations (PDEs). Let $u:\Omega\to\mathbb{R}$ be a function (the state variable) defined on a domain $\Omega\subset\mathbb{R}^{d}$. The governing PDE, together with boundary and initial conditions, can be written compactly as
$\mathcal{F}[u,\theta](\mathbf{x})=\mathbf{0},\qquad\mathbf{x}\in\Omega,$
where $\theta\in\mathbb{R}^{m}$ are the PDE parameters and $\mathcal{F}$ is the associated differential operator, defined by
$\mathcal{F}[u,\theta](\mathbf{x}):=F(\mathcal{D}^{k}u(\mathbf{x}),\dots,\mathcal{D}u(\mathbf{x}),u(\mathbf{x}),\theta),\qquad\mathbf{x}\in\Omega,$ (1)
with $\mathcal{D}^{k}$ denoting the $k$-th order derivative operator. The solution to the PDE depends implicitly on $\theta$; we denote this solution by $u(\cdot,\theta)$.
@@ 2.1 Method Overview
Let $P(\theta)$ be the prior distribution of the PDE parameters. Let $D_{u}$ be the observed data, which is typically a noisy observation of the state $u$ at a set of measurement points. The likelihood $P(D_{u}|\theta)$ measures the probability of observing the data $D_{u}$ given the PDE parameters $\theta$. Since it is usually a functional of the state $u$, thus we also denote it as $\ell[u]$. For example, if we assume $D_{u}=\{\hat{u}_{i}\}_{i=1}^{N}$ with $\hat{u}_{i}=u(\mathbf{x}_{i},\theta)+\eta_{i}$, where $\eta_{i}$ is Gaussian noise of mean 0 and variance $\sigma_{d}^{2}$, then the likelihood of the observed data $D_{u}$ given the PDE parameters $\theta$ is
$\ell[u]=\prod_{i=1}^{N}\frac{1}{\sqrt{2\pi\sigma_{d}^{2}}}\exp\left(-\frac{1}{2\sigma_{d}^{2}}\left|u(\mathbf{x}_{i})-\hat{u}_{i}\right|^{2}\right)$ (2)
@@
which is $P(D_{u}|\theta)$ if $u$ is the solution at $\theta$. Let $P(\theta|D_{u})$ be the posterior distribution of the PDE parameters. By Bayes’ theorem, $P(\theta|D_{u})\propto P(D_{u}|\theta)P(\theta)$. The potential energy, $\mathcal{U}[u,\theta]$, which depends on both the state $u$ and the PDE parameters $\theta$, is defined as the negative of the log-posterior distribution
$\mathcal{U}[u,\theta]=-\log P(\theta|D_{u})=-\log\ell[u]-\log P(\theta)$ (3)
@@
We also denote $U(\theta)=\mathcal{U}[u(\cdot,\theta),\theta]$, as the potential energy ultimately depends on $\theta$ if the state $u$ is the solution at $\theta$. Sampling the posterior distribution with gradient-based MCMCs, such as Langevin Dynamics (LD) or Hamiltonian Monte Carlo (HMC), requires computing the gradient of the potential energy $\nabla_{\theta}U(\theta)$ to effectively explore the posterior distribution.
Bayesian Bilevel Local Operator Learning (B-BiLO).
We consider functions of the form $u(\mathbf{x},\theta)$ that depend on both the spatial-temporal variable $\mathbf{x}$ and the PDE parameters $\theta$. We call $u(\cdot,\theta)$ a local operator at $\theta$ if
$\mathcal{F}[u(\cdot,\theta),\theta](\mathbf{x})=\mathbf{0}\quad(\text{Condition 1}),$
$d_{\theta}\mathcal{F}[u(\cdot,\theta),\theta](\mathbf{x})=\mathbf{0}\quad(\text{Condition 2}),$
for all $x\in\Omega$, where $d_{\theta}$ is the total derivative with respect to $\theta$. Condition 1 requires the residual of the PDE to be zero, i.e., that $u(\cdot,\theta)$ is a solution of the PDE at $\theta$. Condition 2 requires the total derivative of the PDE residual with respect to $\theta$ to be zero, ensuring that the derivative of $u(\cdot,\theta)$ with respect to $\theta$ remains consistent with the PDE constraints. This ensures that small perturbations in $\theta$ lead to small perturbations in $u(\cdot,\theta)$ that still approximately satisfy the PDE constraints. In particular, if we consider a local perturbation $\theta\to\theta+\delta\theta$, a formal Taylor expansion of $\mathcal{F}$ about $\theta$ is zero to second order in $\delta\theta$ (e.g., see [zhangBiLOBilevelLocal2025a]).
@@
In this work, we represent the local operator by a neural network $u(\mathbf{x},\theta;W)$, where $W$ are the weights of the neural network. The two conditions for $u$ to be a local operator at $\theta$ lead to the definition of the local operator loss $\mathcal{L}_{\text{LO}}$:
$\mathcal{L}_{\text{LO}}(\theta,W)=\norm{\mathcal{F}[u(\cdot,\theta;W),\theta]}^{2}_{2}+w_{\rm rgrad}\norm{d_{\theta}\mathcal{F}[u(\cdot,\theta;W),\theta]}^{2}_{2},$ (4)
@@
We call the first term the residual loss and the second term the residual-gradient loss. Both terms can be approximated using the Mean Squared Error (MSE) at a set of collocation points. The weight $w_{\rm rgrad}>0$ is a scalar hyperparameter that controls the relative importance of the residual-gradient loss. In principle, both terms can be minimized to nearly zero. For fixed $\theta$, by minimizing the local operator loss $\mathcal{L}_{\text{LO}}(\theta,W)$ with respect to the weights $W$, we approximate the local operator $u(\mathbf{x},\theta)$ around $\theta$.
@@
We consider the following Bayesian inference problem with a bilevel structure:
[…]
@@
At the upper level, we sample the PDE parameters $\theta$ from the posterior distribution using Hamiltonian Monte Carlo (HMC) [nealMCMCUsingHamiltonian2011]. At the lower level, we train a network to approximate the local operator by minimizing the local operator loss with respect to the weights of the neural network.
Algorithm and Theoretical Analysis
We use HMC with the leapfrog integrator to sample the posterior distribution at the upper level. HMC requires computing the gradient of the potential energy $\nabla_{\theta}U(\theta)$ at various values of $\theta$, which require solving the lower-level problem to some tolerance $\epsilon$. Detailed algorithms are provided in Sec. 4.2. Under mild stability/smoothness assumptions, we show that the error in the gradient $\nabla_{\theta}U(\theta)$ and the error in the posterior distribution, both due to inexact minimization of the lower level problem, are $O(\epsilon)$. The proofs are provided in Sec. 4.3 and SM.
Advantages over existing methods
Many existing methods, such as Bayesian PINNs (BPINNs) [yangBPINNsBayesianPhysicsinformed2021, linkaBayesianPhysicsInformed2022], represent the PDE solution using a Bayesian neural network $v(\mathbf{x};W)$, (note that $\theta$ is not an input to the network), where $W$ follows a prior distribution $P(W)$. Uncertainty in $W$ induces uncertainty in the PDE residual $D_{f}=\{(\mathbf{x}_{f}^{(i)},\eta_{f})\}_{i=1}^{N_{f}}$, $\eta_{f}\sim\mathcal{N}(0,\sigma_{f}^{2})$, and give rise to a likelihood function for the residual term:
$P(D_{f}|W,\theta)=\prod_{i=1}^{N_{f}}\frac{1}{\sqrt{2\pi\sigma_{f}^{2}}}\exp\left(-\frac{1}{2\sigma_{f}^{2}}\left|\mathcal{F}[v(\mathbf{x}_{f}^{(i)};W),\theta]\right|^{2}\right)$ (6)
which serves as a soft constraint to enforce the PDE, and depends on $\sigma_{f}$. Assuming $\theta$ and $W$ are independent, the following joint posterior is sampled using HMC.
These would require sampling the challenging high-dimensional joint posterior of the PDE parameters $\theta$ and network weights $W$. Making $\sigma_{f}$ small enforces the PDE strongly, but also makes the loss ill-conditioned and leads to inefficient sampling. In contrast, B-BiLO avoids these issues by optimizing $W$ deterministically for each $\theta$, and thus only the low-dimensional parameter space $\theta$ is sampled. This approach removes the need for priors on weights, avoids the difficulties setting an artificial residual noise parameter, and enables more stable and efficient exploration of the posterior. A detailed review of BPINNs and it’s challenges are provided in SM.
Efficient Fine-Tuning via LoRA
To increase efficiency, we apply Low-Rank Adaptation (LoRA) [huLoRALowRankAdaptation2021] to reduce the number of trainable parameters at the lower level. Let a hidden-layer weight be $W\in\mathbb{R}^{p\times p}$. Across parameters $\theta$ and an initial guess $\theta_{0}$, BiLO updates only a low-rank increment
$W^{*}(\theta)\approx W^{*}(\theta_{0})+AB,\qquad A\in\mathbb{R}^{p\times r},\;B\in\mathbb{R}^{r\times p},\;r\ll p,$
while keeping the $W^{*}(\theta_{0})$ fixed. We refer to this procedure as “LoRA(rr)”, denoting LoRA fine-tuning with rank $r$ (Algorithm 3), whereas “Full FT” denotes full fine-tuning of all weights (Algorithm 2). Details of the neural network architecture and implementation of LoRA are provided in Methods. 4.1 and 4.4.
Schematic of B-BiLO with LoRA
@@
A schematic of the B-BiLO framework is shown in Figure 1. The figure is based on a model boundary value problem $\theta u_{xx}=\sin(\pi x)$ with the boundary condition $u(0)=u(1)=0$. In this case, the full operator is given by $u(x,\theta)=\sin(\pi x)/(\pi^{2}\theta)$, which solves the PDE for all $\theta>0$. Figure 1 panel (A) illustrates the upper-level problem: the goal is to sample the PDE parameters $\theta$ from the posterior distribution. For gradient-based MCMC methods, we need to compute the gradient of the potential energy $\nabla_{\theta}U(\theta)$ at various values of $\theta$ (orange arrows). Therefore, at any fixed $\theta$, we need access to the local behavior of the potential energy, shown as the blue strip near $\theta_{1}$ and $\theta_{2}$.
@@
In Figure 1 panel (B), the gray surface is the full PDE operator $u(x,\theta)$. At some particular $\theta_{1}$, the local operator $u(x,\theta_{1})$ approximates a small neighborhood of the full operator near $\theta_{1}$ (blue surface). The condition of the local operator is that the residual and the gradient of the residual with respect to $\theta$ are both zero at any fixed $\theta$. Knowing the PDE solution in the neighborhood of $\theta_{1}$ allows us to compute the gradient of the potential energy at $\theta_{1}$.